OpenStudy (christian_10_):

WILL MEDAL!!!!!!! What is the value of the expression? [(3 + 5) • 2 – 5] • 3

OpenStudy (anonymous):

use Bodmas

OpenStudy (anonymous):

so we start with the inner bracket

OpenStudy (anonymous):

(3+5)

OpenStudy (anonymous):

=8*2=16

OpenStudy (anonymous):

now 16-5=11

OpenStudy (anonymous):

now 11*3=33 .. thats all

OpenStudy (christian_10_):

okay, yeah makes since now, I'm not good at math its my worst subject, Lol thank you.

OpenStudy (anonymous):

n.p

OpenStudy (anonymous):

any more questions on this?

OpenStudy (christian_10_):

nope, it was a quiz I got an 80% on it, thanks for all your help! 19/20